OpenStudy (sweetshiva9963):

Write each of the following (base-10) integers in base 2, base 4, and base 8. a) 137 b) 6243 c) 12,345

OpenStudy (loser66):

137 in base 2 is 10001001. if you want to study I will guide you (just one example) and you can find out the leftover.

OpenStudy (whpalmer4):

In base 10, we have the following place values (in reverse order): 1, 10, 100, 1,000, 10,000, etc. In base 2, instead of powers of 10, we have powers of 2: \[1, 2, 4, 8, 16, 32, 64, 128, 256, 512, 1024, 2048, 4096, 8192\] In base 4, we have powers of 4: \[1, 4, 16, 64, 256, 1024, 4096, 16384\] In base 8, we have powers of 8: \[1, 8, 64, 512, 4096, 32768\] To convert, we need to find the combination of powers of the base that equals our number. For example, with 137 written in base 2, the largest power of 2 we can extract is \(2^7=128\). \(137-128=9\), so the next largest power of \(2\) we can extract is \(2^3=8\). \(9-8=1\), so the final power of \(2\) we can extract is \(2^0=1\). We write \(2^7+2^3+2^0\) as \(10001001_2\). For another example, I'll write \(1000_{10} = x_8\) \(1000-512 = 488\) so \(8^3\) is our first power \(488-64 > 64\) so we can take out multiple \(8^2\) \(488 \text{ div } 64 = 7\) so we can take out \(7*8^2\) leaving \(488-448=40\) \(40-8 > 8\) so we can take out multiple \(8^1\), \(5\) in all, leaving \(40-40=0\) So \(1000_{10}=1750_{8}\)